What is the ADR for a four day period given the following information?

Day 1: Revenue=$60,000; Rooms sold=200

Day 2: Revenue=$96,000; Rooms sold=300

Day 3: Revenue =$70,000; Room sold = 250

Day 4: Revenue =$75,000; Rooms sold = 250

a. $301.00

b. $310.00

c. $280.00

d. $320.00
Answer #1

PLEASE RATE POSITIVE IF YOU LIKE MY ANSWER TO PROTECT MY EMPLOYMENT

Total revenue = $ 60,000 + 96000+70000+75000 = $301,000

Total rooms sold = 200+300+250+250 = 1,000

Therefore ADR = $301,000/1,000 = $301

Hence option(a) is the correct answer